OpenStudy (anonymous):

Helpppp If a line has a slope of m=-2/3, find the slope of a line parallel & a line perpendicular to it. and If a line has a slope of m=1/2, find the slope of a line parallel & a line perpendicular to it.

OpenStudy (anonymous):

parallel lines always have same slope for two perpendicular lines, the multiple of their slope is -1 so if one line has slope m, and a line perpendicular to it has slope n then m*n=-1

OpenStudy (anonymous):

that didn't really help with perpendicular part like i understand the parallel but thats it

OpenStudy (anonymous):

ok , so its given that a line has a slope -2/3 suppose the line perpendicular to it has slope x so (-2/3)*x=-1 x =3/2 will be the slope of the other line

OpenStudy (anonymous):

is x always going to = -1?

OpenStudy (anonymous):

no when you multiply the slopes of the two lines, that are perpendicular , you get -1